OpenStudy (anonymous):

find all pairs of integers whose difference is 5 and product is 234. So, what I got is -18 and 13. If i am right how do i go back to check if my answer is right.

OpenStudy (anonymous):

\[x-y=5 \] \[xy=234\] If i go back to check for example:\[-18\times13=-234\]

OpenStudy (freckles):

if you answer is right then the difference of -18 and 13 is 5 (is that a true statement? And also the product of -18 and 13 is 234 (is that a true statement?

OpenStudy (freckles):

both of those statements have to be true.

OpenStudy (freckles):

are they

OpenStudy (anonymous):

the difference of -18 and 13 is not 5 nor the product of -18X13 is not 234. They both give me a negative result

OpenStudy (freckles):

so that means that something went wrong when finding the other number like I believe you got this far (y+18)(y-13)=0 where y=-18 or y=13 right but x=5+y

OpenStudy (freckles):

so if y=-18 then x=5+y=?

OpenStudy (freckles):

and if x=13 then x=5+y=?

OpenStudy (anonymous):

right, i factored it out and thats how i got -18 and 13. so when I go back to check I use the x=5+y equation. So i said x=5+(-18) which x=-13

OpenStudy (freckles):

ok so (-13,-18) is one pair

OpenStudy (freckles):

what about the other one?

OpenStudy (freckles):

i made a type-0

OpenStudy (campbell_st):

well actually there are 2 sets of solutions the numbers are -18 and -13 as well as the positive 18 and 13

OpenStudy (freckles):

and if y=13 then x=5+y=? *

OpenStudy (freckles):

so that means that something went wrong when finding the other number like I believe you got this far (y+18)(y-13)=0 where y=-18 or y=13 right but x=5+y remember we had two y's so we will have two different x's

OpenStudy (freckles):

therefore two different pairs

OpenStudy (anonymous):

ohhh i see my mistake I only solved for y. I didnt solve for x!

OpenStudy (freckles):

well you solved for y which we got was y=-18 or y=13 and we know x=5+y so you already plugged in y=-18 giving you x=5+(-18)=-13 but you also need to plug in the other y which was y=13 into x=5+y

OpenStudy (anonymous):

so x=5+13=18 So y=-18 and 13 and x=18 and -13

OpenStudy (freckles):

well i would list it as actually pairs 18,13 is one pair -18,-13 is another pair

OpenStudy (anonymous):

ah I see, got it! Thank you so much!

OpenStudy (freckles):

if you want list them in order 18,13 -13,-18

OpenStudy (freckles):

I put the bigger number first in each case

OpenStudy (freckles):

so we can have left number-right number=5 in each case

OpenStudy (anonymous):

makes a lot sense now and what was my mistake.

OpenStudy (freckles):

what do you mean

OpenStudy (anonymous):

i meant i get what was my mistake. I only solved for y and i didnt find ALL integers to give me the difference of 5 and the product of two integers to give me 234

OpenStudy (freckles):

in a way i guess you can say it wasn't really wrong it was just incomplete but also -18 and 13 isn't a pair together that satisfies the difference being 5 and the product being 234.

OpenStudy (anonymous):

right! thank you!